Help me solve this equation
Alex scored 13 points in the basketball game. This was 1/5 of the total points the team scored. Write and solve an equation to determine the total points the team scored.

Answer by checkley71(8403)

13=X/5 CROSS MULTIPLY
X=13*5
X=65 POINTS SCORED BY THE TEAM.
PROOF
1/5*65=13
65/5=13
13=13